# DiffPO: Turbocharging Inference Time Alignment of Large Language Models with Diffusion-Style Preference Optimization

### v1

@@ Abstract
Inference-time alignment provides an efficient alternative for aligning LLMs with humans. However, these approaches still face challenges, such as limited scalability due to policy-specific value functions and latency during the inference phase. In this paper, we propose a novel approach, Diffusion-styled Preference Optimization (DiffPO), which provides an efficient and policy-agnostic solution for aligning LLMs with humans. By directly performing alignment at sentence level, DiffPO avoids the time latency associated with token-level generation. Designed as a plug-and-play module, DiffPO can be seamlessly integrated with various base models to enhance their alignment. Extensive experiments on AlpacaEval 2, MT-bench, and HH-RLHF demonstrate that DiffPO achieves superior alignment performance across various settings, achieving a favorable trade-off between alignment quality and inference-time latency. Furthermore, DiffPO demonstrates model-agnostic scalability, significantly improving the performance of large models such as Llama-3-70B. Our model and code are available here.
DiffPO: Diffusion-styled Preference Optimization for Efficient Inference-Time Alignment of Large Language Models
Ruizhe Chen 1,2   Wenhao Chai 3   Zhifei Yang 4   Xiaotian Zhang 1 Joey Tianyi Zhou 5Tony Quek 2   Soujanya Poria 2Zuozhu Liu 1
1 Zhejiang University  2 SUTD  3 University of Washington 4 Peking University  5 A*STAR Centre for Frontier AI Research
@@ 1 Introduction
The alignment of large language models (LLMs) with human preferences has recently emerged as a focal area of research Wang et al. (2023); Shen et al. (2023). Prominent techniques such as Reinforcement Learning from Human Feedback (RLHF) Ouyang et al. (2022) and Direct Preference Optimization (DPO) Rafailov et al. (2024) have demonstrated substantial efficacy. However, these methods require the optimization of individual policies, posing challenges such as high consumption of training resources. Inference-time alignment Mudgal et al. (2023); Han et al. (2024) provides an efficient alternative through direct adjustment of the model’s output distribution, thus avoiding the need for resource-intensive retraining. Despite its advantages, this approach still requires policy-specific value functions, limiting its scalability across different models. Additionally, the inference-time latency remains high, presenting further challenges to its practical deployment.
@@
In this paper, we investigate an efficient and policy-agnostic preference optimization method. We begin by reconsidering the objective of aligning with humans Yao et al. (2023); Shen et al. (2023). As illustrated in Fig. 2(a), the alignment process operates at the sentence level, focusing on adjusting key components of the generated content, such as style or format, to better reflect human intentions or values. Inspired by the global controllability of the diffusion process Li et al. (2022); Lyu et al. (2023), we propose Diffusion-styled Preference Optimization (DiffPO). DiffPO draws an analogy from the diffusion-based denoising process to model the iterative adjustment required for aligning human preferences, as shown in Fig. 2(b). By employing parallel decoding Santilli et al. (2023); Leviathan et al. (2023), DiffPO directly predicts sentence-level transitions, thus avoiding the time latency associated with token-level generation. During the training phase, we optimize the DiffPO with an objective that maps generations with varying alignment levels to an aligned target, making it a policy-agnostic, plug-and-play module. The optimized DiffPO can then be seamlessly integrated with the output of the base model, enhancing its alignment level, as demonstrated in Fig. 2(c).
We evaluate the performance of DiffPO on several benchmark datasets, including AlpacaEval 2 Dubois et al. (2024), MT-bench Zheng et al. (2023), and HH-RLHF Bai et al. (2022). Empirical results demonstrate that DiffPO achieves superior alignment performance across various base models and settings. Compared to inference-time alignment techniques, DiffPO strikes an optimal trade-off between alignment performance and inference-time latency, as shown in Fig. 1. Additional experiments highlight the model-agnostic scalability of DiffPO across different base models. Specifically, DiffPO-9B significantly enhances the performance of models such as Llama-3-70B and GPT-4o, showcasing its capability to improve weak-to-strong supervision.
@@ 2.1 Preliminaries: Large Language Models
During the alignment of LLMs, the objective is to optimize a language model $\pi_{\theta}$ that maximizes the user’s preference Christiano et al. (2017); Ouyang et al. (2022); Rafailov et al. (2024):
@@
In comparison to next-token prediction, where token-level generation is performed sequentially to obtain a sentence, parallel decoding has demonstrated the capacity by enabling sentence-level generation and improving content quality Santilli et al. (2023); Leviathan et al. (2023). Concretely, supposing
@@ 2.2 Diffusion-styled Preference Optimization
The goal of LLM alignment is to align the outputs of LLMs with human values or intentions Yao et al. (2023). In this process, preferences are defined at the sentence-level, focusing on the style or format of complete generated answers, as illustrated in Fig. 2(a). However, the generation of these responses occurs at the token level, following the next-token prediction pattern inherent in LLM modeling. This requires existing alignment techniques to optimize preferences (or rewards) at the token-level, which complicates the learning process Andrychowicz et al. (2017); Zhong et al. (2024); Zeng et al. (2024). This inconsistency prompts us to reconsider the formulation of the alignment process.
@@
Inspired by the potential benefits of the diffusion process in controllable text generation Gong et al. (2022); Han et al. (2022); Ye et al. (2024b), we draw an analogy between the aligning LLMs and the diffusion process. Specifically, we propose Diffusion-styled Preference Optimization (DiffPO), which reconceptualizes alignment as a sentence-level denoising process. The denoising process $\pi$ gradually refines the initial unaligned output $\mathbf{y}^{(0)}$ by adjusting the format or style as a whole. This process ultimately produces the aligned output $\mathbf{y}^{(T)}$, as illustrated in Fig. 2(b). The sentence-level alignment process can be formulated as follows:
@@ 2.3 Consistency Optimization of DiffPO
Inspired by Consistency LLMs Kou et al. (2024), we propose to consistently map any intermediate (unaligned) generation $\mathbf{y}^{(t)}$ to the aligned generation $\mathbf{y}^{(T)}$. We jointly optimize the DiffPO model $\pi_{\theta}$ with two losses: one aligns the intermediate generation with the aligned generation, and the other prevents the corruption of the autoregressive (AR) modeling in the base model, thereby maintaining the generation quality.
@@ 2.4 The Objective of DiffPO within RLHF
In this section, we analyze the role of DiffPO in achieving the goal of RLHF. We start with the same RL objective as prior work, Eq. 2.1, under a general reward function $r^{*}$. Following prior work Peng et al. (2019); Rafailov et al. (2024), the optimal solution to the KL-constrained reward maximization objective in Eq. 2.1 takes the form: $r^{*}(\mathbf{x},\mathbf{y})=\beta\log\left(\frac{\pi^{*}(\mathbf{y}\mid\mathbf{x})}{\pi_{\mathrm{ref}}(\mathbf{y}\mid\mathbf{x})}\right)+\beta\log Z(\mathbf{x})$, where $Z(\mathbf{x})=\sum_{\mathbf{y}}\pi_{\mathrm{ref}}(\mathbf{y}|\mathbf{x})\exp\left(\frac{1}{\beta}r^{*}(\mathbf{x},\mathbf{y})\right)$ is the partition function. With Bradley-Terry model, we can represent the preference function as the difference of rewards for a preferred answer $\mathbf{y}_{w}$ and a dispreferred answer $\mathbf{y}_{l}$:
@@ 2.5 Practical Implementations
To implement DiffPO, we collect the alignment trajectory for each prompt, thereby forming an original training set $\mathcal{D}$. Specifically, for each prompt $\mathbf{x}$ from the UltraFeedback dataset Cui et al. (2023), we generate $T$ responses using different base models. We then employ ArmoRM Wang et al. (2024) reward model to score these responses. The response with the highest score is selected as $\mathbf{y}^{(T)}$. The remaining five responses are ranked based on their scores to form $\mathbf{y}^{(0:T-1)}$. $T$ is set to 6.
@@
During the training phase, we initialize our aligning model $\pi_{\theta}$ using three backbones of varying sizes: Gemma-2-it-2B/9B, and Llama-3-8B-Instruct. The DiffPO model is optimized adhering to the optimization loss in Eq. 6 with parameters $N=256$ and $w=10^{3}$. Given the variable lengths of generations in $\mathcal{D}$, we standardize their lengths through padding or truncation. In the inference phase, the optimized model $\pi_{\theta}^{*}$ is employed to align responses from the vanilla generations produced by base models. Appendix B.1 shows more implementation details.
3 Experiment
@@
3.1 Experiment Setup
Evaluation Benchmarks and Metrics.
We conduct our experiments using two widely recognized benchmarks for open-ended instruction-following: MT-Bench Zheng et al. (2023) and AlpacaEval 2 Dubois et al. (2024). These benchmarks are designed to evaluate the conversational abilities of models across a diverse set of queries. AlpacaEval 2 includes 805 questions drawn from five distinct datasets, while MT-Bench covers eight categories and comprises a total of 80 questions. Additionally, we employ the HH-RLHF Bai et al. (2022) datasets to assess how well the models’ generative capabilities align with human values, particularly emphasizing helpfulness and harmlessness. We adhere to each benchmark’s specific evaluation protocol to report scores. In AlpacaEval 2, we report both the raw win rate (WR) and the length-controlled win rate (LC), comparing performance against the GPT-4 model. In contrast, we present the average score for MT-Bench, also utilizing GPT-4 as the judge model. For HH-RLHF, we report scores that reflect the models’ helpfulness and harmlessness, as well as the overall score. These scores are measured using ArmoRM Wang et al. (2024), a state-of-the-art reward model from RewardBench Lambert et al. (2024), designed to align with human preferences.
@@ 3.1 Experiment Setup
We compare DiffPO with two primary categories of offline preference optimization methods. In the category of training-based methods: Direct Preference Optimization (DPO) Rafailov et al. (2024) reparameterizes reward functions to simplify and stabilize the preference learning process. SimPO Meng et al. (2024) utilizes the average log probability of a sequence as an implicit reward, aligning more closely with model generation. For training-free methods: Black-Box Prompt Optimization (BPO) Cheng et al. (2024) adapts user prompts to better align with LLMs’ input comprehension, achieving user intents optimally without altering LLM parameters. ARGS Khanov et al. (2024) integrates alignment into the decoding process through reward-guided search, eliminating the need for costly RL training. Best-of-N sampling (BoN) Nakano et al. (2021) samples N times and selects the highest-scoring sample based on the reward model, with N set to 4 in our experiments using ArmoRM Wang et al. (2024) as the reward model. Furthermore, Aligner Ji et al. (2024) and MetaAligner Yang et al. (2024a) employ an additional model to learn corrective residuals between preferred and dispreferred responses to refine model generation.
@@
We perform preference optimization primarily on two model families: Llama-3-8B AI@Meta (2024) and Mistral-7B Jiang et al. (2023), under two configurations: SFT and Instruct. In the SFT configuration, we utilize open-source models from SimPO Meng et al. (2024) that follow Zephyr Tunstall et al. (2023) to train the base models (i.e., meta-llama/Meta-Llama-3-8B) on the UltraChat-200k Ding et al. (2023) dataset to derive an SFT model. For the Instruct configuration, we employ an off-the-shelf instruction-tuned models (i.e., meta-llama/Meta-Llama-3-8B-Instruct). To further validate scalability, we conduct additional experiments using the Llama-3.2 series, Qwen-2.5 series Team (2024), and GPT-4o Achiam et al. (2023) as the base models.
@@ 4 Conclusion
Generation Length 256 512 1,024 2,048
Llama-3-SFT 6.21 6.61 6.76 6.71
w. DiffPO-8B ($\Delta$) +0.75 +0.81 +0.93 +1.05
w. DiffPO-9B ($\Delta$) +1.24 +1.64 +1.48 +0.50
Llama-3-Instruct 6.78 7.87 7.99 8.00
w. DiffPO-8B ($\Delta$) +0.24 -0.12 +0.01 +0.02
w. DiffPO-9B ($\Delta$) +0.62 +0.68 +0.34 +0.62
Mistral-SFT 5.73 6.42 6.50 6.36
w. DiffPO-8B ($\Delta$) +1.14 +1.16 +1.25 +1.51
w. DiffPO-9B ($\Delta$) +1.40 +1.47 +1.71 +1.86
Mistral-Instruct 6.39 7.47 7.68 7.64
w. DiffPO-8B ($\Delta$) +0.65 +0.06 +0.13 +0.17
w. DiffPO-9B ($\Delta$) +0.94 +0.59 +0.58 +0.78
Loss MT AE2 HH
GPT-4 LC (%) Avg.
Vanilla Decoding
Llama-3-Instruct 6.78 36.83 0.7985
$L_{\mathrm{AR}}$ 6.90 36.35 0.7971
Llama-3-it w. DiffPO-8B
$L_{\mathrm{AR}}$ only 6.75 35.84 0.7968
$10\times L_{\mathrm{Con}}+L_{\mathrm{AR}}$ 6.85 35.96 0.7997
$100\times L_{\mathrm{Con}}+L_{\mathrm{AR}}$ 6.86 35.92 0.7998
$1,000\times L_{\mathrm{Con}}+L_{\mathrm{AR}}$ 7.02 36.44 0.7998
@@
Appendix A Related Works
A.1 Align LLM with Human Preference.
A prominent approach to learning from human preferences is RLHF Ouyang et al. (2022); Stiennon et al. (2020); Christiano et al. (2017); Bai et al. (2022). In this framework, a reward model is first trained, followed by the training of a bandit policy using Proximal Policy Optimization (PPO) Schulman et al. (2017). Recent advancements such as direct preference optimization (DPO) Rafailov et al. (2024); Meng et al. (2024); Ethayarajh et al. (2024) optimize the bandit policy directly from human preferences, bypassing the need for a reward model. These approaches are simpler to implement and require fewer computational resources. Inference-time approaches, on the other hand, achieve alignment by customizing the output of large language models (LLMs) during the decoding phase, without the need for parameter optimization. This results in enhanced flexibility and efficiency Khanov et al. (2024); Mudgal et al. (2023). One representative method treats the text-generation process as a search problem, guided by external rewards Huang et al. (2024); Han et al. (2024); Chakraborty et al. (2024). Another category of methods focuses on learning to refine the generated text Li et al. (2023a); Ji et al. (2024); Yang et al. (2024a).

### v2

# DiffPO: Diffusion-styled Preference Optimization for Efficient Inference-Time Alignment of Large Language Models

@@ Abstract
Inference-time alignment provides an efficient alternative for aligning LLMs with humans. However, these approaches still face challenges, such as limited scalability due to policy-specific value functions and latency during the inference phase. In this paper, we propose a novel approach, Diffusion-styled Preference Optimization (DiffPO), which provides an efficient and policy-agnostic solution for aligning LLMs with humans. By directly performing alignment at sentence level, DiffPO avoids the time latency associated with token-level generation. Designed as a plug-and-play module, DiffPO can be seamlessly integrated with various base models to enhance their alignment. Extensive experiments on AlpacaEval 2, MT-bench, and HH-RLHF demonstrate that DiffPO achieves superior alignment performance across various settings, achieving a favorable trade-off between alignment quality and inference-time latency. Furthermore, DiffPO demonstrates model-agnostic scalability, significantly improving the performance of large models such as Llama-3-70B.
@@ 1 Introduction
The alignment of large language models (LLMs) with human preferences has recently emerged as a focal area of research [62, 53]. Prominent techniques such as Reinforcement Learning from Human Feedback (RLHF) [47] and Direct Preference Optimization (DPO) [50] have demonstrated substantial efficacy. However, these methods require the optimization of individual policies, posing challenges such as high consumption of training resources. Inference-time alignment [45, 27] provides an efficient alternative through direct adjustment of the model’s output distribution, thus avoiding the need for resource-intensive retraining. Despite its advantages, this approach still requires policy-specific value functions, limiting its scalability across different models. Additionally, the inference-time latency remains high, presenting further challenges to its practical deployment.
@@
In this paper, we investigate an efficient and policy-agnostic preference optimization method. We begin by reconsidering the objective of aligning with humans [65, 53]. As illustrated in Fig. 2(a), the alignment process operates at the sentence level, focusing on adjusting key components of the generated content, such as style or format, to better reflect human intentions or values. Inspired by the global controllability of the diffusion process [38, 43], we propose Diffusion-styled Preference Optimization (DiffPO). DiffPO draws an analogy from the diffusion-based denoising process to model the iterative adjustment required for aligning human preferences, as shown in Fig. 2(b). By employing parallel decoding [51, 37], DiffPO directly predicts sentence-level transitions, thus avoiding the time latency associated with token-level generation. During the training phase, we optimize the DiffPO with an objective that maps generations with varying alignment levels to an aligned target, making it a policy-agnostic, plug-and-play module. The optimized DiffPO can then be seamlessly integrated with the output of the base model, enhancing its alignment level, as demonstrated in Fig. 2(c).
We evaluate the performance of DiffPO on several benchmark datasets, including AlpacaEval 2 [21], MT-bench [74], and HH-RLHF [5]. Empirical results demonstrate that DiffPO achieves superior alignment performance across various base models and settings. Compared to inference-time alignment techniques, DiffPO strikes an optimal trade-off between alignment performance and inference-time latency, as shown in Fig. 1. Additional experiments highlight the model-agnostic scalability of DiffPO across different base models. Specifically, DiffPO-9B significantly enhances the performance of models such as Llama-3-70B and GPT-4o, showcasing its capability to improve weak-to-strong supervision.
@@ 2.1 Preliminaries: Large Language Models
During the alignment of LLMs, the objective is to optimize a language model $\pi_{\theta}$ that maximizes the user’s preference [15, 47, 50]:
@@
In comparison to next-token prediction, where token-level generation is performed sequentially to obtain a sentence, parallel decoding has demonstrated the capacity by enabling sentence-level generation and improving content quality [51, 37]. Concretely, supposing
@@ 2.2 Diffusion-styled Preference Optimization
The goal of LLM alignment is to align the outputs of LLMs with human values or intentions [65]. In this process, preferences are defined at the sentence-level, focusing on the style or format of complete generated answers, as illustrated in Fig. 2(a). However, the generation of these responses occurs at the token level, following the next-token prediction pattern inherent in LLM modeling. This requires existing alignment techniques to optimize preferences (or rewards) at the token-level, which complicates the learning process [4, 75, 69]. This inconsistency prompts us to reconsider the formulation of the alignment process.
@@
Inspired by the potential benefits of the diffusion process in controllable text generation [24, 28, 67], we draw an analogy between the aligning LLMs and the diffusion process. Specifically, we propose Diffusion-styled Preference Optimization (DiffPO), which reconceptualizes alignment as a sentence-level denoising process. The denoising process $\pi$ gradually refines the initial unaligned output $\mathbf{y}^{(0)}$ by adjusting the format or style as a whole. This process ultimately produces the aligned output $\mathbf{y}^{(T)}$, as illustrated in Fig. 2(b). The sentence-level alignment process can be formulated as follows:
@@ 2.3 Consistency Optimization of DiffPO
Inspired by Consistency LLMs [35], we propose to consistently map any intermediate (unaligned) generation $\mathbf{y}^{(t)}$ to the aligned generation $\mathbf{y}^{(T)}$. We jointly optimize the DiffPO model $\pi_{\theta}$ with two losses: one aligns the intermediate generation with the aligned generation, and the other prevents the corruption of the autoregressive (AR) modeling in the base model, thereby maintaining the generation quality.
@@ 2.4 The Objective of DiffPO within RLHF
In this section, we analyze the role of DiffPO in achieving the goal of RLHF. We start with the same RL objective as prior work, Eq. 2.1, under a general reward function $r^{*}$. Following prior work [49, 50], the optimal solution to the KL-constrained reward maximization objective in Eq. 2.1 takes the form: $r^{*}(\mathbf{x},\mathbf{y})=\beta\log\left(\frac{\pi^{*}(\mathbf{y}\mid\mathbf{x})}{\pi_{\mathrm{ref}}(\mathbf{y}\mid\mathbf{x})}\right)+\beta\log Z(\mathbf{x})$, where $Z(\mathbf{x})=\sum_{\mathbf{y}}\pi_{\mathrm{ref}}(\mathbf{y}|\mathbf{x})\exp\left(\frac{1}{\beta}r^{*}(\mathbf{x},\mathbf{y})\right)$ is the partition function. With Bradley-Terry model, we can represent the preference function as the difference of rewards for a preferred answer $\mathbf{y}_{w}$ and a dispreferred answer $\mathbf{y}_{l}$:
@@ 2.5 Practical Implementations
To implement DiffPO, we collect the alignment trajectory for each prompt, thereby forming an original training set $\mathcal{D}$. Specifically, for each prompt $\mathbf{x}$ from the UltraFeedback dataset [16], we generate $T$ responses using different base models. We then employ ArmoRM [61] reward model to score these responses. The response with the highest score is selected as $\mathbf{y}^{(T)}$. The remaining five responses are ranked based on their scores to form $\mathbf{y}^{(0:T-1)}$. $T$ is set to 6.
@@
During the training phase, we initialize our aligning model $\pi_{\theta}$ using three backbones of varying sizes: Gemma-2-it-2B/9B, and Llama-3-8B-Instruct. The DiffPO model is optimized adhering to the optimization loss in Eq. 6 with parameters $N=256$ and $w=10^{3}$. Given the variable lengths of generations in $\mathcal{D}$, we standardize their lengths through padding or truncation. In the inference phase, the optimized model $\pi_{\theta}^{*}$ is employed to align responses from the vanilla generations produced by base models. Appendix 6.1 shows more implementation details.
@@
3 Experiment
3.1 Experiment Setup
Evaluation Benchmarks and Metrics.
We conduct our experiments using two widely recognized benchmarks for open-ended instruction-following: MT-Bench [74] and AlpacaEval 2 [21]. These benchmarks are designed to evaluate the conversational abilities of models across a diverse set of queries. AlpacaEval 2 includes 805 questions drawn from five distinct datasets, while MT-Bench covers eight categories and comprises a total of 80 questions. Additionally, we employ the HH-RLHF [5] datasets to assess how well the models’ generative capabilities align with human values, particularly emphasizing helpfulness and harmlessness. We adhere to each benchmark’s specific evaluation protocol to report scores. In AlpacaEval 2, we report both the raw win rate (WR) and the length-controlled win rate (LC), comparing performance against the GPT-4 model. In contrast, we present the average score for MT-Bench, also utilizing GPT-4 as the judge model. For HH-RLHF, we report scores that reflect the models’ helpfulness and harmlessness, as well as the overall score. These scores are measured using ArmoRM [61], a state-of-the-art reward model from RewardBench [36], designed to align with human preferences.
@@ 3.1 Experiment Setup
We compare DiffPO with two primary categories of offline preference optimization methods. In the category of training-based methods: Direct Preference Optimization (DPO) [50] reparameterizes reward functions to simplify and stabilize the preference learning process. SimPO [44] utilizes the average log probability of a sequence as an implicit reward, aligning more closely with model generation. For training-free methods: Black-Box Prompt Optimization (BPO) [14] adapts user prompts to better align with LLMs’ input comprehension, achieving user intents optimally without altering LLM parameters. ARGS [34] integrates alignment into the decoding process through reward-guided search, eliminating the need for costly RL training. Best-of-N sampling (BoN) [46] samples N times and selects the highest-scoring sample based on the reward model, with N set to 4 in our experiments using ArmoRM [61] as the reward model. Furthermore, Aligner [31] and MetaAligner [63] employ an additional model to learn corrective residuals between preferred and dispreferred responses to refine model generation.
@@
We perform preference optimization primarily on two model families: Llama-3-8B [3] and Mistral-7B [32], under two configurations: SFT and Instruct. In the SFT configuration, we utilize open-source models from SimPO [44] that follow Zephyr [60] to train the base models (i.e., meta-llama/Meta-Llama-3-8B) on the UltraChat-200k [20] dataset to derive an SFT model. For the Instruct configuration, we employ an off-the-shelf instruction-tuned models (i.e., meta-llama/Meta-Llama-3-8B-Instruct). To further validate scalability, we conduct additional experiments using the Llama-3.2 series, Qwen-2.5 series [59], and GPT-4o [1] as the base models.
@@ 3.2 Experiment Results
Generation Length 256 512 1,024 2,048
Llama-3-SFT 6.21 6.61 6.76 6.71
w. DiffPO-8B ($\Delta$) +0.75 +0.81 +0.93 +1.05
w. DiffPO-9B ($\Delta$) +1.24 +1.64 +1.48 +0.50
Llama-3-Instruct 6.78 7.87 7.99 8.00
w. DiffPO-8B ($\Delta$) +0.24 -0.12 +0.01 +0.02
w. DiffPO-9B ($\Delta$) +0.62 +0.68 +0.34 +0.62
Mistral-SFT 5.73 6.42 6.50 6.36
w. DiffPO-8B ($\Delta$) +1.14 +1.16 +1.25 +1.51
w. DiffPO-9B ($\Delta$) +1.40 +1.47 +1.71 +1.86
Mistral-Instruct 6.39 7.47 7.68 7.64
w. DiffPO-8B ($\Delta$) +0.65 +0.06 +0.13 +0.17
w. DiffPO-9B ($\Delta$) +0.94 +0.59 +0.58 +0.78
Table 5: Loss and Hyperparameter Ablation. The results indicate that applying DiffPO-9B to the base model (i.e., gemma-2-9B) yields outperforming performance than vanilla decoding of any single models, demonstrating the effectiveness of DiffPO strategy.
Loss MT AE2 HH
GPT-4 LC (%) Avg.
Vanilla Decoding
Gemma-2-9B-it 7.37 54.15 0.8265
$L_{\mathrm{AR}}$ 7.42 58.19 0.8279
Gemma-2-9B-it w. DiffPO-9B
$L_{\mathrm{AR}}$ only 7.41 58.01 0.8301
$10\times L_{\mathrm{Con}}+L_{\mathrm{AR}}$ 7.56 59.27 0.8389
$100\times L_{\mathrm{Con}}+L_{\mathrm{AR}}$ 7.55 59.43 0.8435
$1,000\times L_{\mathrm{Con}}+L_{\mathrm{AR}}$ 7.60 60.56 0.8438
@@ 4 Conclusion
Supplementary Material
The supplementary material is structured as follows:
  • Related works in Section 5.
  • Additional experiment setup and results in Section 6.1.
  • Additional analysis and ablation studies in Section 7.
5 Related Works
5.1 Align LLM with Human Preference.
Large Language Models (LLMs) shows impressive performance in many tasks [54, 71, 55, 72, 9]. A prominent approach to learning from human preferences is RLHF [47, 58, 15, 5, 13]. In this framework, a reward model is first trained, followed by the training of a bandit policy using Proximal Policy Optimization (PPO) [52]. Recent advancements such as direct preference optimization (DPO) [50, 44, 22, 73] optimize the bandit policy directly from human preferences, bypassing the need for a reward model. These approaches are simpler to implement and require fewer computational resources. Inference-time approaches, on the other hand, achieve alignment by customizing the output of large language models (LLMs) during the decoding phase, without the need for parameter optimization. This results in enhanced flexibility and efficiency [34, 45]. One representative method treats the text-generation process as a search problem, guided by external rewards [30, 27, 10]. Another category of methods focuses on learning to refine the generated text [39, 31, 63].